OpenStudy (anonymous):

Solve the equation. –6 + 3x = –9 A. –1 B. –6 C. –5 D. –3

OpenStudy (baldymcgee6):

Move the -6 over to the right hand side.

OpenStudy (anonymous):

Isolate the variable. \[ -6+3x=-9\\ 3x=-9+6=-3\\ x=\frac{-3}{3}=-1 \]